OpenStudy (anonymous):

an angle whose supplement is 54 degrees less than 4 times it's complement. Find the measure of given angle

OpenStudy (anonymous):

Complement = c = 90 - x Supplement = s = 180 -x s = 4c -54 180-x = 360 - 4x - 54 3x = 126 x = 42